Derive the equation (23) at page 42 of Unit 2, which represents the polar equation
of a conic when the directrix L corresponding to a focus F is taken to the right of F.
1
Expert's answer
2020-03-10T10:41:39-0400

Let "F" be a fixed point, the focus, and let "L" be a fixed line, the directrix, in a plane. A conic section, or conic, is the set of all points "P" in the plane such that 


"{PF \over PL}=e, e>0"

where "e" is a fixed positive number, called the eccentricity.

If "e=1," the conic is a parabola.

If "e<1," the conic is an ellipse.

If "e>1," the conic is a hyperbola.

By locating a focus at the pole, all conics can be represented by similar equations in the polar coordinate system. In each of these equations,

"(r, \theta)" is a point on the graph of the conic.

"e" is the eccentricity. (Remember that "e>0").

"p" is the distance between the focus (located at the pole) and the directrix.

For a conic with a focus at the origin, if the directrix is "x=\pm p," where "p" is a positive real number, and the eccentricity is a positive real number "e," the conic has a polar equation


"r={e\cdot p \over 1\pm\cos{\theta}}"

For a conic with a focus at the origin, if the directrix is "y=\pm p,"where "p" is a positive real number, and the eccentricity is a positive real number "e," the conic has a polar equation


"r={e\cdot p \over 1\pm \sin{\theta}}"

Given that the directrix "L" corresponding to a focus "F" is taken to the right of "F."

Directrix "L: x=p"



The distance from the focus to the point "P" in polar is just "r."


"PF=r"

The distance from the point "P" to the directrix "x=p" is "PL=p-r\cos{\theta}." Then


"e={PF \over PL}={r \over p-r\cos{\theta}}"

Solve for "r"


"e\cdot p-e\cos{\theta}\cdot r=r""r={e\cdot p \over 1+ \cos{\theta}}"

The conic has a polar equation


"r={e\cdot p \over 1+ \cos{\theta}}"